A patient with aspiration pneumonia has an order for aminophylline 250 mg in 500 mL of D5W to infuse between 0.5 and 0.7 mg/kg/hr. The patient weighs 82.4 kg.

a. What is the mg/hr dose range? ___________________________________________

b. What is mL/hr flow rate? ____________
Answer #1

Aminophylline 0.5 to 0.7 mg/kg/hr = 0.5 \times 82.4 = 41.2 mg/hr

                                                     = 0.7 \times 82.4 = 57.68 or 57.7 mg/hr

a. The mg/hr dose range is 41.2 to 57.7 mg/hr

b. The required medication = required dose \times available volume \div available dose

                                            = 41.2 \times 500 \div 250

                                            = 82.4 ml

The required medication = required dose \times available volume \div available dose

                                       = 57.7 \times 500 \div 250

                                      = 115.4 ml

The flow rate range is 82.4 ml/hr to 115.4 ml/hr